OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5). A. 23 B. –53 C. –9 D. 3

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

Sub g(x) for x in F(x) then sub -5 for x in (f ° g)

OpenStudy (anonymous):

i dont know what this is

OpenStudy (anonymous):

What is g(x)?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what is f(x)

OpenStudy (anonymous):

i dont know this is on my test

OpenStudy (anonymous):

that is timed

OpenStudy (anonymous):

i have an hour to answer 20 questions

OpenStudy (anonymous):

Substitute -4x+3 for x in (-2x-7)

OpenStudy (anonymous):

i dont know how to substitute

OpenStudy (anonymous):

(f ° g)(-5) = -2g(-5) - 7 = -2(-4(-5) + 3) - 7 = -53

OpenStudy (anonymous):

-2(-4x+3)-7

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

the answer you get you then substitute -5 for x

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

multiply -2(-4x+3)

OpenStudy (anonymous):

what do you mean -13? from my previous post: (f ° g)(-5) = -2g(-5) - 7 = -2(-4(-5) + 3) - 7 = -2(20 +3) -7 = -2(23) -7 = -46 -7 = -53 can't get more straight forward than this.....

OpenStudy (anonymous):

@sirm3d

OpenStudy (sirm3d):

substitute x=-5 in g(x)

hero (hero):

f(x) = -2x - 7 g(x) = -4x + 3. f o g (x) = f(g(x)) f(g(x)) = -2(-4x + 3) - 7 f(g(x)) = 8x - 6 - 7 f(g(x)) = 8x - 13 f(g(-5)) = 8(-5) - 13 f(g(-5)) = -40 - 13 f(g(-5)) = -53

OpenStudy (sirm3d):

hahaha, too much algebra @Hero

hero (hero):

If you want, I can include descriptions of each step

OpenStudy (sirm3d):

this is simpler. substitute -5 in g(x), and the result you get, substitute in f(x)

hero (hero):

I just showed that my steps are simpler. The simpler result is f(g(x)) = 8x - 13 You reduce f(x) and g(x) to a simpler form, f(g(x)). Now you plug in f(g(-5) f(g(-5)) = 8(-5) - 13 f(g(-5)) = -40 - 13 = -53

OpenStudy (anonymous):

im about to go somewhere is it A

hero (hero):

No it's not A, which is very sad

OpenStudy (sirm3d):

it is A for the first substitution @jennyjewell25 . use that value of f(x) to get the correct answer

hero (hero):

Trust me @sirm3d, you saying "it is A" probably confused her.

OpenStudy (anonymous):

what lol

OpenStudy (sirm3d):

no i did not. i said A is what you'll get in the first substitution

OpenStudy (sirm3d):

@jennyjewell25 when you use -5 on g(x), you get 23. Now use that value on f(x) and you'll get the correct answer.

OpenStudy (anonymous):

i dont get it its not i have to go

hero (hero):

I told you she's confused @sirm3d. She thinks 23 is the answer to the whole thing now.

hero (hero):

I knew she was confused the moment she asked if 23 was the answer right after I posted the full set of steps.

OpenStudy (sirm3d):

g(x)=-4x+3 g(-5) = 23

OpenStudy (sirm3d):

f(x) = -2x-7

OpenStudy (sirm3d):

f(23) = ??

hero (hero):

@sirm3d, if you want to help, you're gonna have to learn how to post things in a step by step manner. If you only post g(x)=-4x+3 g(-5) = 23 Then a user may not still understand the process because you skipped a crucial intermediary step. Now if you post this: g(x)=-4x+3 g(-5) = -4(-5) + 3 g(-5) = 20 + 3 g(-5) = 23 They might have a better chance of understanding the process.

OpenStudy (sirm3d):

@Hero we'll, if you want her to learn, teach her how to fish, not serve her fish.

hero (hero):

My friend, even the fisherman needed to know how to fish as one is not born knowing how. Part of learning is seeing the process done. When the person was young they were not a fisherman yet. They watched mom, pop, and uncle joe do it. Once they were of age, then mom and pop guided them. Combining what they already observed along with the guidance enabled them to finally understand. But the observation was a critical part of that process. The importance of observation should never be undermined.

OpenStudy (sirm3d):

@Hero you have to assume that the topic has been taught before taking the test. i'm sure the fisherman was shown how to catch fish. it's her time to fish.

hero (hero):

My friend, entire civilizations, countries, companies, organizations, and individuals were destroyed because they decided to "make an assumption" before starting a project only to realize that making that assumption was the wrong approach and thus at that point, it was too late to reverse the process. Making assumptions is exactly the wrong approach. Never make assumptions. Always make sure you understand the situation fully before embarking on the journey.

hero (hero):

The victims of Hurricane Sandy made an assumption. They assumed everything would be okay because they have been through huge storms in the past. But they assumed wrong and now many have died.

OpenStudy (sirm3d):

then you assumed that they assumed wrong.

OpenStudy (anonymous):

I think the jenneyjewell25 appreciates all the help given to her (I think), so let move on and help other people, but on another note, I believe the correct answer is the one I gave previously which is -53 which is the same as Hero's answer.

hero (hero):

It's called covering all bases. Since you cannot assume what someone knows or what they don't know, you provide the complete set of steps, that way, if they already know something, then they can just skip to the steps they don't know. But allow them the opportunity to do that. If you skip steps when presenting the answer, then you will have skipped the steps they don't understand. Then they will still be completely clueless. So yes, I will assume they don't know because it is the best and only assumption to make.

hero (hero):

Not everyone has a classroom or teacher to run to. For some of them Open Study is their only hope.

OpenStudy (sirm3d):

what is wrong if i let her do the substitution? i did not say that 23 is the correct answer to the problem. i said it is the result of the first substitution, and i put there that there is another substitution coming.

hero (hero):

You can let her do the substitution, but only if she knows how to do it. If she doesn't know how to do it, then she will need to see the process done. In this case, she clearly did not. Thus, the only other option is to show her the process.

hero (hero):

I usually guide students when I know for sure that they already know how to do certain steps. If I don't know what they can do, I provide the full set of steps. That's how I operate. That's why I have 99 smart score.

hero (hero):

In your case, you might want to ask the student questions before deciding to help or guide them.

hero (hero):

Because believe it or not, you can actually end up confusing a student more, which would be counter-productive.

hero (hero):

@sirm3d